OpenStudy (anonymous):

Assume that there are 9 board members: 6 females, and 3 males including Huck. There are 4 tasks to be assigned randomly, including that of notifying members of meeting times. (1) Find the probability that Huck is given a task. (2) Find the probability that Huck is given the task of notifying members of meeting times.

OpenStudy (kropot72):

(1) Assume that only one task is given to each member. The number of ways of assigning the 4 tasks with Huck included in each is 8C3. The total number of ways of assigning the four tasks among all the members is 9C4. Therefore the probability of Huck getting a task is \[\frac{8C3}{9C4}=\frac{4}{9}\]

OpenStudy (kropot72):

(2) In each combination of 4 tasks there is a 1/4 probability of Huck's getting the specified task. Therefore the probability that Huck is chosen for a task and is assigned the specified task is \[\frac{4}{9} \times \frac{1}{4}\]

OpenStudy (anonymous):

Thanks kropot!

OpenStudy (kropot72):

You're welcome :)